OpenStudy (anonymous):

If the scale factor of two similar solids is 3 : 11, what is the ratio of their corresponding areas? What is the ratio of their corresponding volumes?

OpenStudy (mathmate):

Areas are proportional to the square of the ratios, and volumes, cubes. Example: scale factor: 2:5 ratio of areas: 2^2:5^2 =4:25 ratio of volumes : 2^3 : 5^3 = 8 : 125 This is because for areas, we multiply together two dimensional units, and for volumes, three.

OpenStudy (mayankdevnani):

Recall that if the linear dimensions of two similar solids are in ratio a:b, then their surface areas are in ratio a^2:b^2 and their volumes are in ratio a^3:b^3... Since the side lengths of two similar solids are in ratio 3:7, their volumes are in ratio 3^3:11^3 = 27:1331..... ok @sue101

OpenStudy (anonymous):

ok thanks!!